A = P(1+r)t (t is as an exponent)
In the formula A = P(1+r)t, P is the principal\, r is the annual rate of interest, and A is the amount after t years. An account earning interest at a rate of 4% has a principal of $500,000. If no more deposits or withdrawls are made, about how much money will be in the account after five years?
A $705,200
B $620,700
C $608,300
D $575,000

Answer by ankor@dixie-net.com(22740)

A = P(1+r)t (t is as an exponent)
In the formula A = P(1+r)t, P is the principal\, r is the annual rate of interest, and A is the amount after t years. An account earning interest at a rate of 4% has a principal of $500,000. If no more deposits or withdrawls are made, about how much money will be in the account after five years?
:
A = 500000(1.04)^5
A = 500000*1.2166529
A = $608,326.45, so guess it's c